Pt 2 what is the profit margin?
QUESTION 43 1 points Balance Sheet Cash &Ms Receivables Inventories ts Payable 80,000 Notes Payable Fixed Assets otal Assets Annual Sales$1,500,000 ies Common Equity 280,000 ot Cost of Goods Sold-$1.200,000 Number of Shares Outstanding 20 000 Net Income $60,000 What is the days sales outstanding (DSO) of the firm? (Use a 360-day year.) O 12.0 days O 13.2 days 16.5 days O 22 days O 30.5 days Click Save and Submit to sawe and submit. Click Save All Answers to save all answers
Answer #1

1)

Receivables turnover = Net credit sales / average receivables

Receivables turnover = 1,500,000 / 50,000

Receivables turnover = 30

DSO = 360 / 30

DSO = 12.0 days

2)

Profit margin = ( net income / sales ) * 100

Profit margin = ( 60,000 / 1,500,000) * 100

Profit margin = 4.0%